OpenStudy (anonymous):

An Integral a Day: day 3 \[\int_{-\infty}^\infty e^{ax^2}dx\]

OpenStudy (turingtest):

this will only converge if a<0

OpenStudy (turingtest):

double integral in polar coordinates?

OpenStudy (turingtest):

oh, I don't know that... is that like the CPV ?

OpenStudy (experimentx):

not really sure ... values are usually in pi

OpenStudy (experimentx):

and i would love to derive that.

OpenStudy (turingtest):

I got up to\[\int_{-\infty}^{\infty}e^{-ax^2}dx=2\int_{0}^{\infty}e^{-ax^2}dx;a>0\]\[u=ax^2\implies x=\sqrt{\frac ua}\]\[du=2axdx\implies dx=\frac{du}{2\sqrt{au}}\]and the integral becomes\[\frac1{\sqrt a}\int_{0}^{\infty}u^{-1/2}e^{-u}du\]at which point I'm stuck.

OpenStudy (turingtest):

apparently (according to wolfram) the integral above converges to \(\sqrt\pi\) but I have no idea how to show that

OpenStudy (experimentx):

looks like gamma function

OpenStudy (turingtest):

not something I've worked with, though I sure know about it

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Gamma_function it sure is a gamma function.

OpenStudy (turingtest):

ah, and it specifically mentions the proof for the gamma function of argument 1/2 is \(sqrt\pi\)

OpenStudy (turingtest):

\(\sqrt\pi\)*

OpenStudy (experimentx):

1/a * integ 0 to inf e^-u t^(1-3/2) du

OpenStudy (experimentx):

that is equal to 1/sqrt(a) * ( gamma(3/2) = 1/2 * sqrt(pi)) that is our answer.

OpenStudy (turingtest):

yep :) cool, first time I'm understanding it so we have\[\frac1{\sqrt a}\Gamma(\frac12)\]

OpenStudy (experimentx):

you made mistake .. gamma 3/2

OpenStudy (experimentx):

must be 1-z

OpenStudy (turingtest):

yes, my mistake like I said, I'm new to the gamma function

OpenStudy (experimentx):

i think it got something to do with real analysis.

OpenStudy (experimentx):

check this out too ... it occurs frequently has some kinda relation with gamma function ... though i forgot. http://en.wikipedia.org/wiki/Beta_function

OpenStudy (turingtest):

wait, no, my exponent is negative 1-z=1-1/2=-1/2 so its gamma(1/2) isn't it?

OpenStudy (turingtest):

that should read z-1=1/2-1=-1/2 dyslexia!

OpenStudy (experimentx):

no ,,, you had -1/2 to begin with and that is 1-3/2

OpenStudy (turingtest):

yes, but \[\Gamma(z)=\int_{0}^{\infty}e^{-u}u^{z-1}\]plug in z=1/2 and you get my integral

OpenStudy (turingtest):

I'm sorry if I'm confused, I just want to understand...

OpenStudy (experimentx):

now i am confused.

OpenStudy (turingtest):

scroll down in the wiki article to "general principles" it specifically mentions \[\Gamma(\frac12)=\sqrt\pi\]

OpenStudy (experimentx):

1-z = -1/2

OpenStudy (experimentx):

yeah i know that ... and gamma works like factorial.

OpenStudy (turingtest):

but it is defined as t^(z-1)

OpenStudy (turingtest):

not 1-z...

OpenStudy (experimentx):

looks like ... i am nuts. you are right.

OpenStudy (turingtest):

ok, sweet! I finally learned something :)

OpenStudy (turingtest):

thanks for hanging in there with me, I needed to think it through

OpenStudy (experimentx):

lol ... it was the first thing written in wikipedia article.

OpenStudy (turingtest):

momentary dyslexia so frustrating

OpenStudy (experimentx):

;)

OpenStudy (experimentx):

and the answer was concluded with \[ \sqrt{\frac{\pi}{a}}\]

OpenStudy (turingtest):

...with \(a>0\) The original problem has \(e^{ax^2}\)so that answer would be technically be\[\sqrt{\frac\pi {-a}};a<0\]but same thing of course

OpenStudy (anonymous):

Turing, you were right, double integral in polar coordinates is the way I went. Also I made a mistake and missed a minus sign in the exponent writing it down. \[ I = \int_{-\infty}^\infty e^{-ax^2}dx\]\[I^2 = \int_{-\infty}^\infty\int_{-\infty}^\infty e^{-a(x^2+y^2)}dxdy\]then switch to polar:\[I^2 = \int_0^\infty \int_0^{2\pi}\rho e^{-a\rho^2}d\rho d\phi\]\[I^2 = 2\pi \int_0^\infty \rho e^{-a\rho^2}d\rho\]substitution as follows: \[ \rho^2 =u ~~~~~~~~ du = 2\rho d\rho \] and our integral becomes: \[I^2 = \pi \int_0^\infty e^{-a u} du = \frac{-\pi e^{-a u }}{a}|_0^\infty \] which evaluates to: \[I^2 = lim ~~u\to \infty~~( \frac{-\pi e^{-au}}{a} + \frac{\pi}{a}) = \frac{pi}{a}\]\[and~~so~~I = \sqrt{\frac{\pi}{a}}\]